OpenStudy (anonymous):

For f(x) = 5 – 3x and g(x) = x^2 – 1, find f(g(x)) and g(f(x)).

OpenStudy (hoblos):

f(g(x))= 5- 3(x^2 - 1) = -3x^2 +8 g(f(x))= (5 – 3x)^2 -1 = 25 - 30x + 9x^2 -1 = 9x^2 -30x +24

OpenStudy (anonymous):

Thank you. I got the same two answers, and my test was saying -3x^3+8 was the answer.